A body freely falling from a tower covers a distance of S=h/2 in the last second of its motion.the height of tower is?

Grade:11

1 Answers

shivang belwariar
11 years ago

we have the formula - S = U + a/2(2N-1) [ WHERE U = INITIAL SPEED , & N = value of TIME IN Tth SECOND . ]so using the formula ------h/2 = u + a/2(2t-1)h/10 = 2t-1 { u=0 & a=10}t = h+10/20so, total time = h+10/20now by formula = h=ut + 1/2at^2, (u=0 a=10)h=h^2+100+20h/80